Petroleum pollution in ocean stimulates the growth of certain bacteria. An assessment of this group has been made by counting the bacteria in each of the five randomly chosen specimens of ocean water (of a fixed size). The count obtain were as follows 65, 43, 46, 62, 54 find the standard deviation of a sample of numbers round answer to two decimal places.


1
Expert's answer
2021-01-12T00:45:25-0500

Mean "μ = \frac{65+43+46+62+54}{5} = 54"



Standard deviations

"σ = \sqrt{\frac{1}{n-1}\sum^{n}_{i=1}(x_i-μ)^2} \\

= \sqrt{\frac{370}{5-1}} \\

= \sqrt{92.5} \\

= 9.61"

Answer: 9.61